OpenStudy (anonymous):

Semi-major axis length and semi-minor axis length of (x-2)%2 + 16(y-4)^2 = 4

terenzreignz (terenzreignz):

Ellipse?

OpenStudy (anonymous):

Yeah :D

terenzreignz (terenzreignz):

Well, first thing you wanna do with ellipses is arrange them in this fashion... \[\huge \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]Can you do that? We just need the right side to be equal to 1, which it isn't, currently...

OpenStudy (anonymous):

Yeah I did it but I'm not if it's right hehe

terenzreignz (terenzreignz):

Show me what you did, then ;)

OpenStudy (anonymous):

\[\frac{ (x-2)^{2} }{ 4 } + \frac{ (y-4)^{2} }{ \frac{ 1 }{ 4 } } = 1\]

OpenStudy (anonymous):

Gosh that was hard getting that right on here haha

terenzreignz (terenzreignz):

You... have been holding out on me. You're correct :)

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

Okay just needed to check :p

OpenStudy (anonymous):

I have another question... can I type it here?

terenzreignz (terenzreignz):

Now, once you've gotten it in this form \[\huge \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] You need to find a and b, which should be easy enough?

terenzreignz (terenzreignz):

Wait, we're done with this question? LOL What are your answers?

OpenStudy (anonymous):

Yeah I'm done haha, semi-major is 2, semi-minor is 1/2 ?

terenzreignz (terenzreignz):

Nicely done :) Sure, put up your next question~

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Thanks for the medal :D

OpenStudy (anonymous):

Find the Cartesian equation of \[\left| z-2i \right| = \sqrt{2}\left| z-1 \right|\]

terenzreignz (terenzreignz):

That was quite a jump. Hang on.

OpenStudy (anonymous):

Haha yeah... okay thanks :D

terenzreignz (terenzreignz):

What does Cartesian equation even mean? :/

OpenStudy (anonymous):

z = x + yi

terenzreignz (terenzreignz):

I see. Well, first, I'd think you should substitute x+yi for z...

terenzreignz (terenzreignz):

Note: This is the first time I've heard of this, so go easy :D

OpenStudy (anonymous):

Yeah I kinda did it, but it got really confusing and I couldn't continue haha

OpenStudy (anonymous):

Oh okay haha

terenzreignz (terenzreignz):

Well, if we substitute, we should get \[\Large |x+yi-2i|=\sqrt2|x+yi-1|\] right?

OpenStudy (anonymous):

Isn't \[|z-2i|=\sqrt{(z-2i)(z+2i)}\]

OpenStudy (anonymous):

It's complex numbers so I did |x + (y-2)| = root2 |(x-1) + y|

OpenStudy (anonymous):

Okay wait

terenzreignz (terenzreignz):

@agostino it sure looks like it, but z itself is complex. Best separate it into real and imaginary components first. \[\Large |x+(y-2)i|=|(x-1)+yi|\] This is when it's rearragned, @stephanieee._

OpenStudy (anonymous):

|z - 2i| is \[\sqrt{x ^{2} + (y-2)^{2}}\]

OpenStudy (anonymous):

Yeah I got that :D

terenzreignz (terenzreignz):

That it is, @stephanieee._ Are you sure you needed help with this? :/

OpenStudy (anonymous):

Yeah but they want it in cartesian :(

OpenStudy (anonymous):

So I have to expand and simplify I think

terenzreignz (terenzreignz):

That shouldn't be too hard? :)

terenzreignz (terenzreignz):

So you have \[\Large \sqrt{x^2+(y-2)^2}=\sqrt{(x-1)^2+y^2}\]

OpenStudy (anonymous):

The answer I got was \[-\frac{ 40 }{ 9 } = (x-2)^{2} + (y-\frac{ 2 }{ 3 })^{2} \] and that does not look right, I think I'm meant to have 1 on the right hand side

terenzreignz (terenzreignz):

Yikes... I forgot the square root of 2 bit... \[\Large \sqrt{x^2+(y-2)^2}=\sqrt2\sqrt{(x-1)^2+y^2}\]

OpenStudy (anonymous):

Yep and then square both sides?

terenzreignz (terenzreignz):

So, we square both sides...yes \[\large x^2+(y-2)^2=2(x-1)^2+2y^2\]

terenzreignz (terenzreignz):

And let the world tremble... \[\large x^2+y^2-4y+4=2x^2-4x+2+2y^2\]

terenzreignz (terenzreignz):

As it sees all you are about to accomplish... \[\Large 2=x^2-4x+y^2-4y\]

OpenStudy (anonymous):

Oh wait I think I got it now, I stuffed up on one of the calcs! LOL

terenzreignz (terenzreignz):

It still leads to something negative on the left side, though... I think. Is that okay?

OpenStudy (agent0smith):

@terenzreignz since you started with complex numbers, i think that's fine.

OpenStudy (anonymous):

Should be fine, I have to go eat now anyway :D

OpenStudy (anonymous):

Thank you so much!

terenzreignz (terenzreignz):

That's reassuring. I was kinda hoping it'd lead to an equation of an ellipse, though.

OpenStudy (anonymous):

Thanks again :)

OpenStudy (agent0smith):

Oh but weren't you really finding the 'magnitude' of the complex numbers? Which is just a scalar isn't it...

terenzreignz (terenzreignz):

I know... it's all perplexing z = x+yi where x and y are real... so why did this happen?

terenzreignz (terenzreignz):

Must have screwed up with the algebra bit.

OpenStudy (agent0smith):

well i guess x and/or y can still be negative...

OpenStudy (anonymous):

Haha I think I just have to change to cartesian form :/

terenzreignz (terenzreignz):

I got my wish (ellipse) after all, and no ordinary ellipse, but a circle :D

OpenStudy (anonymous):

Yay thanks!! :D

terenzreignz (terenzreignz):

@stephanieee._ WAIT I DID screw up some algebra. Hold your horses...

terenzreignz (terenzreignz):

\[\Large \sqrt{x^2+(y-2)^2}=\sqrt2\sqrt{(x-1)^2+y^2}\] Squaring both sides \[\Large {x^2+(y-2)^2}=2[(x-1)^2+y^2]\] Simplifying \[\Large x^2+y^2-4y+4 = 2[x^2-2x+1+y^2]\]\[\Large x^2+y^2-4y+4 = 2x^2-4x+2+2y^2\] \[\Large 4-2=2x^2-x^2-4x+2y^2-y^2+4y\]\[\Large 2=x^2-4x+y^2+4y\]

OpenStudy (agent0smith):

^yep, i just found that. It looks like it'll change the center of the circle. \[\Large 2=x^2-4x+y^2+4y \]\[\large 2 +4+4 = (x-2)^2 + (y+2)^2 \] \[\large 10 = (x-2)^2 + (y+2)^2 \]

terenzreignz (terenzreignz):

Still a circle though @stephanieee._ Please tell me you saw this -.-

OpenStudy (anonymous):

Yeah I saw it haha

terenzreignz (terenzreignz):

Just one little plus sign. If you were writing it, just one vertical slash through the minus sign in the "y-2" bit ;) Good hunting :D

OpenStudy (anonymous):

Hahah okay :)